Find n to be a perfect square trinomial
x^2+20x+n
I have looked inside of my algebra book in the index and in sections on perfect square trinomials, and I found absolutely no problems like this in my textbook. So please help me on this problem I am just confused. thank you so much for your time, it really means alot to me.

Answer by scott8148(6628)

think about factoring a quadratic

the factors of the constant term (n) have to add to the coefficient of the x term (20)

for the quadratic to be a perfect square, the factors have to be the same
__ if f is a factor, then 2f=20 so f=10
__ and f^2=n=100

x^2+20x+100=(x+10)^2